OpenStudy (anonymous):

The pairs of vectors v(1)=(-1,1,0) and v(2)=(1,0,1) span a plane P in R(3). The pairs of vectors w(1)=(0,1,0) and w(2)=(1,1,0) span a plane Q in R(3). Show that P and Q are different and compute the subspace of R(3) that is given by the intersection P(intersect)Q? @helder_edwin

OpenStudy (anonymous):

If you have some time to help.

OpenStudy (anonymous):

To show that they are diferent you can think like this: if they were equal then the vector product of one of the vectors of one plane by one of the vectors of the other plane, must give a vector that is perpendicular to both planes, and this vector must be paralel to the vector that is defined by the formula of the plane, therefore their vector product must be zero. When you do it, you'll see that it is not zero and then the planes are not the same.

OpenStudy (anonymous):

Or you can find the vector that is given by the formula by doing the vector product of the 2 vectors they already give you.

OpenStudy (anonymous):

Did I manage to explain it? I believe its kind of confusing.

OpenStudy (anonymous):

okay, so if i do the cross product of v1 and v2, then do the cross product of w1 and w2, are you saying that they should be the same or multiplies of each other and that would mean they would be the same plane, and if the cross products are different then they would be different planes? is that what you are saying?

OpenStudy (anonymous):

Yes thats it, if the cross products are multiples of each other then they are the same, and vice versa

OpenStudy (anonymous):

okay... well that solves the first part... now how do you find subspace intersection P(intersect)Q

OpenStudy (anonymous):

You first need to find a point that belongs to both planes, then you take the same cross products, and make their cross products to find a vector thatbelongs to both planes, with a point and a vector, you have the line.

OpenStudy (anonymous):

and how do you find that point?

OpenStudy (anonymous):

@TuringTest do you know how to find this intersection

OpenStudy (anonymous):

uggh i need help rissa lol @Smile4rissa

OpenStudy (anonymous):

I took the cross product of the first pair and got (1,1,-1) and the cross product of the second pair and got (0,0,-1) so that means that since they are different vectors then P and Q are different planes? But I don't understand how to find the intersection?

OpenStudy (anonymous):

i found an example online that took the sum of the two vectors in the span of each and set them equal to each other. so P=span{(1,-1,0), (0,1,1)} .... (x,-x+y,y) Q=span{(1,0,0),(0,1,0)} .... (a,b,0) so setting those two equal x=a -x+y=b y=0 but then i got lost what the example did after this... but I am not sure if this is even right

OpenStudy (helder_edwin):

give a second. i remember there's a very easy way: i'm looking for it.

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

r u familiar with strang's books on linear algebra? if so, this is in one if them.

OpenStudy (anonymous):

no i am not

OpenStudy (helder_edwin):

first find the null space of \[ \large \begin{pmatrix} -1 & 1 &0 & 1\\ 1 & 0 & 1 & 1\\ 0 & 1 & 0 & 0 \end{pmatrix} \]

OpenStudy (helder_edwin):

u should read them (just as reference) they're very good.

OpenStudy (anonymous):

I found another example that took the null space of P-Q. but I wasn't sure about that either.. but I will do that one quick

OpenStudy (anonymous):

okay so I got (-1,0,-2,1)

OpenStudy (helder_edwin):

i got the null space is spanned by (1,0,-2,1)

OpenStudy (anonymous):

ohh yes you are right, sign error on my part.. so is that the intersection?

OpenStudy (helder_edwin):

no.

OpenStudy (anonymous):

so then once we find the null space, how do we use that to find the intersection?

OpenStudy (helder_edwin):

the first three columns have pivots. and the fourth column is the independent variable. right?

OpenStudy (helder_edwin):

\[ \large \color{red}{1}\cdot(-1,1,0)+\color{red}{0}\cdot(1,0,1) \color{red}{-2}\cdot(0,1,0)+\color{red}{1}\cdot(1,1,0)=(0,0,0) \] agree?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

how do I use that to find the intersection?

OpenStudy (helder_edwin):

the first three vectors correspond to the pivots and the last one to the free variable.

OpenStudy (helder_edwin):

so u can write \[ \large 1(-1,1,0)-2(0,1,0)=-1(1,1,0) \] and this vector spans the intersection.

OpenStudy (helder_edwin):

\[ \large P\cap Q=\langle(1,1,0)\rangle \]

OpenStudy (anonymous):

okay so it is just a line then.

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

was that correct above, to take the cross products of each set to see if the planes are different?

OpenStudy (helder_edwin):

well. what u proved with that is that they are not parallel. hence they must intersect. also, remember that P and Q aren't just planes, they r subspaces. so, should u have obtained parallel normal vectors then that would have been a proof that P and Q are equal. This would not work with planes (translated subspaces of dim = 2), though.

OpenStudy (anonymous):

is there a better way to show that P and Q are different?

OpenStudy (helder_edwin):

i don't know if it is better, but you could prove that the vectors that span P don't belong to Q and viceversa. but the hint @ivanmlerner gave you is pretty cool.

OpenStudy (anonymous):

would it matter which vector you took from which plane when doing that?

OpenStudy (anonymous):

@ivanmlerner what you said did kind of confuse me.

OpenStudy (anonymous):

What I sugested was: v1xv2 gives a vector that is perpendicular to P w1xw2 gives a vector that is perpendicular to Q Since all vectors that belongs to P are perpendicular to the first cross product and all the vectors that belong to Q are perpendicular to the second cross product. Because of this when you do the cross product of the cross product: (v1xv2)x(w1xw2) you must get a vector that belongs to both planes. Now, this vector gives a direction to the line that is the intersection of both planes, and all you need now is a point to locate this line in space that belongs to both planes. Then, the formula for the line is: p+ (v1xv2)x(w1xw2) where p is the point.

OpenStudy (anonymous):

Oh, wait, they dont give you any point in P or Q, so they are general planes. Then you cant find the point p but the direction is still ok.

OpenStudy (anonymous):

would I still get (1,1,0) doing it this way?

OpenStudy (anonymous):

Let me see

OpenStudy (anonymous):

b/c I think I got (-1,1,0)

OpenStudy (anonymous):

I got this too

OpenStudy (anonymous):

so was what I did earlier to get (1,1,0) wrong? doing it this way that should be the intersection right?

OpenStudy (anonymous):

Look, I didn't have linar algebre so I don't understand the method above, but I think this one is correct, but it might not, @helder_edwin could you take a look here?

OpenStudy (helder_edwin):

i don't think that \((v_1\times v_2)\times(w_1\times w_2)\) will necesarily span the intersection of P and Q.

OpenStudy (anonymous):

strangs book you mentioned earlier is the same guy that teaches at the MIT?

OpenStudy (anonymous):

I found somewhere that said this \[P(x)=Q(y) so P(x)-Q(y)=0\] and then the matrix would be \[\left[\begin{matrix}-1 & 1 & 0 & -1\\ 1 & 0 & -1 & -1\\ 0 & 1 & 0 & 0\end{matrix}\right]\] and then the Null space=(-1,0,-2,1) but I don't know if this makes sense or not.

OpenStudy (helder_edwin):

u can easily prove that (1,1,0) is both in P and Q. the following systems MUST be consistent: \[ \large \left(\begin{array}{cc|c} -1 & 1 & 1\\ 1 & 0 & 1\\ 0 & 1 & 0 \end{array}\right)\qquad\qquad \left(\begin{array}{cc|c} 0 & 1 & 1\\ 1 & 1 & 1\\ 0 & 0 & 0 \end{array}\right) \]

OpenStudy (helder_edwin):

yes, @ivanmlerner . Strang teaches at MIT.

OpenStudy (anonymous):

that makes sense.

OpenStudy (anonymous):

Then why is my answer wrong? I just want to know.

OpenStudy (helder_edwin):

yes, it is pretty much the same. since u r looking for the intersection of two spaces for which u have bases then \[ \large av_1+bv_2=cw_1+dw_2 \] is the system whose solutions (if they exist) are in both spaces.

OpenStudy (anonymous):

Sorry I have a very limited knowledge of linear algebra, my answer was based on geometry. Could you try to explain to me wy my answer is wrong with geometric arguments?

OpenStudy (helder_edwin):

i think u r right.

OpenStudy (helder_edwin):

the first system is inconsistent

OpenStudy (anonymous):

wait. you think his answer is right?

OpenStudy (helder_edwin):

actually it is.

OpenStudy (helder_edwin):

i'm checking what went wrong with my approach.

OpenStudy (helder_edwin):

sorry. i just can't figure out what went wrong. what i did should have given us the base of the intersection.

OpenStudy (helder_edwin):

the problem with @ivanmlerner 's geometric approach is that it only works in R^3. so u need a more general method.

OpenStudy (anonymous):

Yes, that's true.

OpenStudy (helder_edwin):

@zonazoo . if u leave this post open, i'll try to post tomorrow night the answer using linear algebra.

OpenStudy (anonymous):

I would, but we can still respond if the post is closed right?.. I have two more problems that I am really having trouble with that I want to post.

OpenStudy (anonymous):

i definitely still want to find if there is a more general approach

OpenStudy (anonymous):

@helder_edwin is it not possible to use my mthod with more dimensions by analogy?

OpenStudy (helder_edwin):

yes. it would make sense. but in dimensions higher than 3, we have nothing ressembling the cross product.